10. The Bok Chicken Factory is trying to figure out how to minimize the cost of producing 1200 units of chicken parts. The production function is q = 100L0.5 K0.5. The wage rate is birr 9 per hour and the rental rate on capital is birr 4 per machine hour



A.B. Find the maximum output that can be produced for a total cost of birr 720.

1
Expert's answer
2022-02-03T12:05:09-0500

w=9

r=4

q=1200

"q=100L^\frac{1}{2}K^\frac{1}{2}"

"(\frac{\alpha}{\beta})(\frac{K}{L})= (\frac{w}{r})"


"(\frac{0.5}{0.5})(\frac{K}{L})=(\frac{9}{4})"


"\frac{K}{L}=\frac{9}{4}"

K="\frac{9}4L"

Substitute K in the production function

"q=100(\frac{13}{4})^\frac{1}{2}(L^\frac{1}{2})"


"1200=100(\frac{13}{4}L)^\frac{1}{2}"


"L^\frac{1}{2}=(\frac{1200}{(100)(3.25^\frac{1}2)})"


"L=(\frac{1200}{(100)(3.25^\frac{1}2)})^2"

"= \frac{1440000}{32500}= 44.31"


"K= \frac{9}{4}L"

"K= \frac{9}{4}\times 44.31= 99.69"


TC= wL+rK

720=9L+4K


But L="\frac {9}{4}K"

"720= 9L+\frac{9}{4}L"

"720=\frac{45}{4}L"

"L=\frac{(720\times 4)}{45}= 64"



"K=\frac{9}{4}\times 64= 144"



Total Output=144+64= 208